# The Silicon Ceiling: Auditing GPT’s Race and Gender Biases in Hiring

### v2

@@ 3.3.1. Generating GPT Resumes
For our second study, we used the names selected as described above (Table 1), and asked GPT to generate a resume given only a name, to probe its latent biases. Our prompt reads as follows: “[NAME] is applying for a position in the United States. Write a detailed and concise resume for [NAME]. Only report up to three jobs, and give exact start and end dates. Include the following sections: Education, Skills, and Work Experience.” We included the U.S. stipulation, as well as the request for a “concise” resume, with specific sections to ideally ensure the resumes produced are comparable to each other and limit nationality-related bias in the task.
@@ 4.1. Study 1 - Resume Assessment
Figure 3. Average scores for gender vary based on gender representation for occupations with more than 50% women compared with less than 50% women; in most prompts, men’s names scored higher than women’s. Bars illustrate the 95% confidence intervals.
@@ 4.2. Study 2: Resume Generation
In our second experiment, we prompted GPT to create resumes, 320 given 10 trials of the same 32 names across 2 genders (men and women) and 4 racial and ethnic groups (Black, Hispanic or Latinx, Asian, and White). We then followed the protocol outlined in Section 3.3.2, GPT-generated resumes for the attributes of interest identified through our initial iterative coding: education level, college major, graduation date, work experience, gaps in work, seniority and domain for each job listed, and potential markers of immigrant status (like non-native English, non-English languages, or work experience outside of the U.S.).
4.2.1. No Evidence of Maternity Bias
4.2.2. Gender Biases in Work Experience
Visualizing resumes’ years of experience, shown in Figure 5, we observe that GPT assigned women’s resumes, for all racial and ethnic groups except Asian, less work experience than men’s resumes. Running a linear mixed-effects model with name as a random effect, we confirm these findings: women’s resumes had significantly fewer months of experience ($\beta=-16.260,p=0.022$), as did Asians’ resumes ($\beta=-17.654,p=0.014$), while the intersection of the two showed that Asian women’s resumes had more experience ($\beta=18.387,p=0.060$). For the full test results, see Appendix Table  7.
Resumes were also coded for the seniority in each job listed on each resume. Jobs were coded according to the U.S. Bureau of Labor Statistics Occupational Outlook Handbook [^4] https://www.bls.gov/ooh/, where jobs were given a 0 if they did not require a college degree, 1 if they were internships or jobs that require some college, 2 if they were entry-level positions that required an undergraduate degree, and 3 if they were a senior position or required a graduate degree.
@@ 4.2.4. Stereotypes and Other Biases in Resume Work Experience
Finally, we noticed a recency bias in the data; all resumes generated listed graduation years between 2010-2022, with 72.0% of resumes with graduate dates between 2018-2022. This trend was also observed in the total years of work experience, since most applicants only had a few years (visualized in Appendix Figure 10). We also saw a consistent level of education: all resumes had bachelor degrees, and none had further degrees.
[…]
@@ 5.1. Biases in Resume Assessment
Our results indicate that GPT exhibits biases in both Resume Assessment and Generation. Answering our first research question, When scoring prospective candidates in a hiring setting, does GPT display biases along the lines of race, gender, or their intersection?, we did not find drastic differences in scores based on gender and racial across all prompts, but did find some biases in several prompts. For the Rating and Interview prompts, GPT tended to assign higher scores to men than women, while for the Hiring prompt, the observed trend aligned with gender stereotypes, with women receiving higher scores in majority-woman fields, but not in majority-men ones. We observed similar stereotypical trends in terms of race; GPT scored White names higher in occupations where White people are over-represented.
While the differences between scores by race-gender group are very small, we still found statistically significant differences and considerable variation across several prompts. In terms of future audits, this underscores the importance of investigating a wide range of prompts. A recent study prompted GPT in Dutch for a score of whether to invite job candidates for an interview in the Dutch labor market, finding no gender bias and instead gender-race interactions effects and preference for Dutch names  (Lippens, 2024). Our findings, when prompting for interview scores in English for the U.S. labor market, did not find interaction effects. This suggests that there may be language and market differences in GPT performance, similar to linguistic work on gender bias differences in different languages (Menegatti and Rubini, 2017). Taken together, these findings emphasize the need for culturally-specific and international auditing efforts.
@@ 5.3. Reflecting on Potential Sources of Bias
In our studies, GPT exhibited gender and racial biases, along with others. These observed biases may originate from the training data; prior work has shown that training data for other AEDTs is frequently non-representative, biased, or inaccurate (Ajunwa, 2020a). For instance, many AEDTs are trained largely on men’s resumes, which can lead to models scoring women lower (Ajunwa, 2020b). They are also often trained on top performers and current employees data, which may not be a representative sample (Sánchez-Monedero et al., 2020). LLMs, meanwhile, are trained not on top performance in any domain, but rather with online data; according to OpenAI, GPT models were trained on data from the Common Crawl public dataset of web pages.
As previous works have stated (Ray, 2023), using online data may not truly reflect the state of the real world, and even when it does, may contain biases; we see these in our study. For instance, reflecting on the skew toward younger resumes in the Study 2, we note that it is common for example resumes to be posted online to help young job seekers prepare to transition from college to careers. As a result, examples and templates may skew towards younger applicants. It is also possible that younger applicants are more likely to post their resumes online, due to conducting more of their daily lives online. While we cannot conclude what led to the biases we observed, a fundamental limitation of algorithm auditing, we encourage future work that builds and analyzes LLMs for such biases in training data and elsewhere.
@@ 5.4. Implications for Large Language Models as Hiring Tools
While we do not find large differences in our Assessment study, our results are statistically significant and pervasive. Taken in conjunction with the results of our Resume Generation task, and with prior work on bias in LLMs which has shown that they learn biases that mirror humans’ implicit biases (Caliskan et al., 2017), exhibit racial and gender bias (Abid et al., 2021; Kotek et al., 2023; Ding et al., 2023), have occupational biases (Kirk et al., 2021), and make up rationalizations for bias (Kotek et al., 2023), the need to reconsider their use in hiring is clear. Our findings support prior work (Ajunwa, 2019; Raghavan et al., 2020; Sloane et al., 2022; Sánchez-Monedero et al., 2020) that suggests the use of automated tools in hiring, far from making those processes “fair” or “objective”, may instead perpetuate existing social biases. Following the idea of the glass ceiling, a metaphor for the invisible barrier limiting the advancement of women in the workplace (Cotter et al., 2001), we call this bias against marginalized individuals perpetuated by computing processes the silicon ceiling.
Importantly, race and gender need not be the only axes on which individuals suffer under the silicon ceiling. But like most resume studies, which most commonly focus on race and gender (Bertrand and Mullainathan, 2004; Daniel, 1968; Levinson, 1975; Chen et al., 2018), recent legislation has the same scope. New York City’s Local Law 144, which inspired this project, requires employers using automated hiring tools to report the results of bias audits to ‘prove’ fairness by race, gender, and the intersection of the two. This focus may be too myopic, given that our results suggest age, educational status, and nationality biases may be more significant, at least for this particular system. We encourage future researchers and policymakers to consider a wider range of potential issues when auditing automated systems.
@@ 5.5. Limitations & Future Work
We note a few limitations that also provide opportunities for future study. First, our results are limited by the names and resumes we used. While we did select them based on validated prior work, and controlling for socioeconomic indicators, U.S. citizenship, and English proficiency, our results do hinge directly on a relatively small number of names that do not cover all groups of people. Since the resumes in the first study for each occupation were identical except their names, all biases observed are specifically due to that difference. On real resumes, however, other markers may amplify or interact with those biases, so our study likely underestimates the biases that GPT will exhibit in real-world settings. We encourage future work to expand in scope, studying a wider range of identity categories and with more granularity within each. Second, in the resume assessment task, we observed differences in GPT scores between occupations with fewer marginalized people compared to those with more diversity. However, such results could be caused by some other, unobserved variables — for example, some occupations are more likely to use resumes as part of their application process. We recommend future work continue to explore this space and study a greater number of occupations. Lastly, there is limited research on how LLMs are used in practice, and a need for future work exploring their use in situ (e.g., to see where hiring managers draw the line on interviewing candidates). As models are constantly changing, auditing guidelines  (Metaxa et al., 2021b) call for audits to be repeated regularly, when new versions are released or system updates are suspected.
@@ 6. Conclusion
As large language models (LLMs) are increasingly applied to workplace settings and hiring contexts, bias audits are increasingly necessary to evaluate whether they perpetuate existing biases or introduce new ones. To explore the potential impact of LLMs on hiring practices, we conducted a two-study algorithm audit of race and gender biases in GPT-3.5 taking inspiration from prior work on correspondence and resume audits. In Study 1, we focused on resume assessment, varying the name by gender and race indicators and resumes for different roles and asking GPT to score resumes according using three prompts (overall rating, willingness to interview, and willingness to hire). We found evidence for some scoring differences based on race, race representation, and gender representation across several prompts. Our second study explored resume generation to further probe GPT’s underlying biases related to the names used in Study 1. In this study, we prompted GPT to create resumes based on the same names, finding evidence of gender and racial biases: women’s names resulted in resumes with lower seniority, work experience industry assignment contained stereotypes, and Asian and Hispanic names had more immigrant markers, such as non-native English and non-U.S. education and work experiences, whereas Black and White names did not. When the results of the two studies are taken in conjunction, we find evidence for the model upholding existing gender and racial biases based on stereotypes. Our findings contribute to a growing body of literature on LLM biases, in particular when used in workplace contexts, as well as methodology for conducting algorithm audits of automated employment decision tools.
@@ Appendix A Appendix
Table 3. GPT prompts used in the Resume Assessment task. The Rating prompt only asks for an overall score with no context. The Interview prompt asks for a score on how likely GPT would invite a candidate back for an interview. The Hiring prompt asked for a score on how likely GPT would hire the candidate.
@@
Figure 7. Number of trials vs. mean score for matched software engineer occupation that was used to pick an appropriate number of trials in the Resume Assessment task.
@@
Figure 8. Average scores for Rating, Interview, and Hiring scenario prompts comparing matched and mismatched pairs of resumes and job descriptions, (separated by scores for anonymous resumes and scores for named resumes) show that anonymous resumes perform both higher or lower than the named resumes for different prompts, with no clear pattern. Due to the sporadic nature of the anonymous resumes, we omit them from any further analysis. The 95% confidence intervals are omitted from the figure because they are too small to be meaningfully displayed.
@@
Table 4. Linear mixed effects model of resume scores for Rating Prompt
@@
Table 5. Linear mixed effects model of resume scores for Interview Prompt
@@
Table 6. Linear mixed effects model of resume scores for Hiring Prompt
@@
Table 7. Linear mixed effects model of experience (in months) for resume generation task
@@
Table 8. Linear mixed effects model of seniority for resume generation task
@@
Figure 9. A histogram of gap lengths shows that most resumes did not contain gaps. Nine-month gaps are the next most common, but further analysis shows this to be due to the length of the average school year, with gaps occurring between internships.

### v3

@@ 3.3.1. Generating GPT Resumes
For our second study, we used the names selected as described above (Table 1), and asked GPT to generate a resume given only a name, to probe its latent biases. Our prompt reads as follows: “[NAME] is applying for a position in the United States. Write a detailed and concise resume for [NAME]. Only report up to three jobs, and give exact start and end dates. Include the following sections: Education, Skills, and Work Experience.” We included the U.S. stipulation and request for a “concise” resume with specific sections to ideally ensure the resumes produced are comparable to each other and limit nationality-related bias.
@@ 4.1. Study 1 - Resume Assessment
Figure 3. Average scores for gender vary based on gender representation for occupations with more than 50% women compared to less than 50% women; in most prompts, men’s names scored higher than women’s. Bars illustrate the 95% confidence intervals.
@@ 4.2. Study 2: Resume Generation
In our second experiment, we prompted GPT to create 320 resumes, given 10 trials of the same 32 names, covering 2 genders (men and women) and 4 racial and ethnic groups (Black, Hispanic or Latinx, Asian, and White). We then followed the protocol outlined in Section 3.3.2, GPT-generated resumes for the attributes of interest identified through our initial iterative coding: education level, college major, graduation date, work experience, gaps in work, seniority and domain for each job listed, and potential markers of immigrant status (like non-native English, non-English languages, or work experience outside of the U.S.).
@@
4.2.1. Gender Biases in Work Experience
Visualizing resumes’ years of experience, shown in Figure 5, we observe that GPT assigned women’s resumes, for all racial and ethnic groups except Asian, less work experience than men’s resumes. Running a linear mixed-effects model with name as a random effect, we confirm these findings: women’s resumes had significantly fewer months of experience ($\beta=-16.260,p=0.022$), as did Asians’ resumes ($\beta=-17.654,p=0.014$). For the full test results, see Appendix Table  7.
Resumes were also coded for the seniority in each job listed on each resume. Jobs were coded according to the U.S. Bureau of Labor Statistics Occupational Outlook Handbook, where jobs were given a 0 if they did not require a college degree, 1 if they were internships or jobs that require some college, 2 if they were entry-level positions that required an undergraduate degree, and 3 if they were a senior position or required a graduate degree.
4.2.2. No Evidence of Maternity Bias
@@ 4.2.4. Stereotypes and Other Biases in Resume Work Experience
Finally, we noticed a recency bias in the data; all resumes generated listed graduation years between 2010-2022, with 72.0% of resumes with graduation dates between 2018-2022. This trend was also observed in the total years of work experience, since most applicants only had a few years (visualized in Appendix Figure 10). We also saw a consistent level of education: all resumes had bachelor degrees, and none had further degrees.
@@ 5.1. Biases in Resume Assessment
Our results indicate that GPT exhibits biases in both Resume Assessment and Generation. Answering our first research question, When scoring prospective candidates in a hiring setting, does GPT display biases along the lines of race, gender, or their intersection?, we did not find drastic differences in scores based on gender and race across all prompts, but did find some biases in several prompts. For the Rating and Interview prompts, GPT tended to assign higher scores to men than women, while for the Hiring prompt, the observed trend aligned with gender stereotypes, with women receiving higher scores in majority-woman fields, but not in majority-men ones. We observed similar stereotypical trends in terms of race; GPT scored White names higher in occupations where White people are over-represented.
While the differences between scores by race-gender group are very small, we still found statistically significant differences and considerable variation across several prompts. In terms of future audits, this underscores the importance of investigating a wide range of prompts. A recent study prompted GPT in Dutch for a score of whether to invite job candidates for an interview in the Dutch labor market, finding no gender bias and instead gender-race interaction effects and preference for Dutch names  (Lippens, 2024). Our findings, when prompting for interview scores in English for the U.S. labor market showed different effects. This suggests that there may be language and market differences in GPT performance, similar to linguistic work on gender bias differences in different languages (Menegatti and Rubini, 2017). Taken together, these findings emphasize the need for culturally-specific and international auditing efforts.
@@ 5.3. Reflecting on Potential Sources of Bias
In our studies, GPT exhibited gender and racial biases, along with others. These observed biases may originate from the training data; prior work has shown that training data for other AEDTs is frequently non-representative, biased, or inaccurate (Ajunwa, 2021). For instance, many AEDTs are trained largely on men’s resumes, which can lead to models scoring women lower (Ajunwa, 2020b). They are also often trained on top performers and current employees data, which may not be a representative sample (Sánchez-Monedero et al., 2020). LLMs, meanwhile, are trained not on top performance in any domain, but rather with online data; according to OpenAI, GPT models were trained on data from the Common Crawl public dataset of web pages.
As previous works have stated (Ray, 2023), using online data may not truly reflect the state of the real world, and even when it does, may contain biases; we see these in our study. For instance, reflecting on the skew toward younger resumes in Study 2, we note that it is common for example resumes to be posted online to help young job seekers prepare for the transition from college to careers. As a result, examples and templates may skew towards younger applicants. It is also possible that younger applicants are more likely to post their resumes online, due to conducting more of their daily lives online. But we cannot draw definitive conclusions about the root causes of the the biases we observed, and we encourage future work to investigate LLMs for such biases in training data and elsewhere.
@@ 5.4. Implications for Large Language Models as Hiring Tools
While we do not find large differences in our Assessment study, our results are statistically significant and pervasive. Taken in conjunction with the results of our Resume Generation task, and with prior work on bias in LLMs which has shown that they learn biases that mirror humans’ implicit biases (Caliskan et al., 2017), exhibit racial and gender bias (Abid et al., 2021; Kotek et al., 2023; Ding et al., 2023), have occupational biases (Kirk et al., 2021), and make up rationalizations for bias (Kotek et al., 2023), the need to reconsider their use in hiring is clear. Our findings support prior work (Ajunwa, 2020a; Raghavan et al., 2020; Sloane et al., 2022; Sánchez-Monedero et al., 2020) that suggests the use of automated tools in hiring, far from making those processes “fair” or “objective,” may instead perpetuate existing social biases. Following the idea of the glass ceiling, a metaphor for the invisible barrier limiting the advancement of women in the workplace (Cotter et al., 2001), we call this bias against marginalized individuals perpetuated by computing processes the silicon ceiling.
Importantly, race and gender need not be the only axes on which individuals suffer under the silicon ceiling. But like most resume studies, which most commonly focus on race and gender (Bertrand and Mullainathan, 2004; Daniel, 1968; Levinson, 2014; Chen et al., 2018), recent legislation has the same scope. New York City’s Local Law 144, which inspired this project, requires employers using automated hiring tools to report the results of bias audits to ‘prove’ fairness by race, gender, and the intersection of the two. This focus may be too myopic, given that our results suggest age, educational status, and nationality biases may be more significant, at least for this particular system. We encourage future researchers and policymakers to consider a wider range of potential issues when auditing automated systems.
@@ 5.5. Limitations & Future Work
We note a few limitations that also provide opportunities for future study. First, our results are limited by the names and resumes we used. While we did select them based on validated prior work, and controlling for socioeconomic indicators, U.S. citizenship, and English proficiency, our results do hinge directly on a relatively small number of names that do not cover all groups of people. Since the resumes in the first study for each occupation were identical except their names, all biases observed are specifically due to that difference. On real resumes, however, other markers may amplify or interact with those biases, so our study likely underestimates the biases that GPT will exhibit in real-world settings. We encourage future work to expand in scope, studying a wider range of identity categories and with more granularity within each. Second, in the resume assessment task, we observed differences in GPT scores between occupations with fewer marginalized people compared to those with more diversity. However, such results could be caused by some other, unobserved variables — for example, some occupations are more likely to use resumes as part of their application process. We recommend future work continue to explore this space and study more occupations. Lastly, there is limited research on how LLMs are used in practice, and a need for future work exploring their use in situ (e.g., to see where hiring managers draw the line on interviewing candidates). As models are constantly changing, auditing guidelines  (Metaxa et al., 2021b) call for audits to be repeated regularly, when new versions are released or system updates are suspected.
@@ 6. Conclusion
As large language models (LLMs) are increasingly applied to workplace settings and hiring contexts, bias audits are increasingly necessary to evaluate whether they perpetuate existing biases or introduce new ones. To explore the potential impact of LLMs on hiring practices, we conducted a two-study AI audit of race and gender biases in GPT-3.5 taking inspiration from prior work on correspondence and resume audits. In Study 1, we focused on resume assessment, varying the name by gender and race indicators and resumes for different roles and asking GPT to score resumes according using three prompts (overall rating, willingness to interview, and willingness to hire). We found evidence for some scoring differences based on race, race representation, and gender representation across several prompts. Our second study explored resume generation to further probe GPT’s underlying biases related to the names used in Study 1. Prompting GPT to create resumes based on the same names, we found evidence of gender and racial biases: women’s names resulted in resumes with less work experience, and the industries assigned reflected social stereotypes. Additionally, Asian and Hispanic names presented markers of immigration status, such as non-native English skills and non-U.S. education and work experiences, while Black and White names did not. Taking these studies’ results in conjunction, we find evidence for latent gender and racial biases based on stereotypes. Our findings contribute to a growing body of literature on LLM biases, in particular when used in workplace contexts, as well as methodology for conducting AI audits of automated employment decision tools.
7. Positionality & Ethical Considerations
Our research team represents a range of racial, gender, and other identities, and is primarily based in the United States. This is reflected by the framing of this work, which focused on the United States hiring context, though we situate it in international research on the topic, and encourage further global explorations.
To conduct this work ethically, we took care to design the study to have the best chance of identifying biases where they exist, to examine bias intersectionally and broadly, and to ensure correctness in our method’s execution. We made sure, for example, to analyze for interaction effects of race and gender, and in Study 2, which involved qualitative analysis, to consider a wide range of potential biases, including by maternity status, nationality/immigration status, and age. We also selected names with care to control for confounding variables, and asked Career Services professionals to review the resume materials used.
A potential adverse impact of this work would be its use in justifying the application of LLMs to hiring tasks because the magnitudes of the biases we find in Study 1 are small. We make it clear throughout the paper that we do not condone such practices, nor believe them to be supported by our findings. As Study 2 confirms, the model reflects reveal latent biases, including ones we did not test for in Study 1. Work in this area is still developing, and we expect more evidence to continue emerging and confirming that tasks with the potential to impact people’s livelihoods ought to be made in validated, transparent, and accountable ways.
Acknowledgements.
We thank and appreciate Leon Yin and Davey Alba, technology reporters and data journalists at Bloomberg, for their inspiring reporting and generous collaboration in forming our research question and designing Study 1.
@@ Appendix A Appendix
Table 3. GPT prompts included a Rating, Interview, and Hiring scenario for the Resume Assessment task.
@@
Figure 7. Number of trials vs. mean score for the software engineer occupation was used to pick an appropriate number of trials for the Resume Assessment task.
@@
Figure 8. Average scores separated by matched and mismatched pairs of resumes and job descriptions for named and anonymous resumes show that anonymous resumes perform both higher or lower than the named resumes for different prompts with no clear pattern. Due to the sporadic nature of the anonymous resumes, we omit them from any further analysis. The 95% confidence intervals are omitted because they are too small to be meaningfully displayed.
@@
Table 4. Linear mixed effects model of resume scores for Rating Prompt.
@@
Table 5. Linear mixed effects model of resume scores for Interview Prompt.
@@
Table 6. Linear mixed effects model of resume scores for Hiring Prompt.
@@
Table 7. Linear mixed effects model of experience (in months) for resume generation task.
@@
Table 8. Linear mixed effects model of seniority for resume generation task.
@@
Figure 9. A histogram of gap lengths shows that most resumes did not contain gaps. Nine-month gaps are the next most common, but further analysis shows it is due internships with gaps occurring during the length of the average school year.